OpenStudy (cedric):

Assume a 125 pound person burns 150 calories after 15 minutes of jogging and 200 calories after 20 minutes of jogging. Part 1 (2 points): Write an equation to represent this data. Use calories as the y-coordinate and minutes as the x-coordinate. Part 2 (2 points): What is the slope of the equation found in part 1? What does the slope represent in terms of the information provided? Part 3 (2 points): How many calories will this person burn after 35 minutes of jogging? Using complete sentences, explain how the equation, slope, or graph can help to predict calories burned.

OpenStudy (anonymous):

(15, 150) and (20, 200); rise over run, (200-150) / (20 - 15) = 50 / 5 = a slope of 10 (you should be able to see this right away in that the relationship between the two is obviously that the latter is 10 times the former. y = mx + b; we know y, m, and x at the first ordinate 150 = 10*15 + b. b is zero, so y = 10x + 0 (you don't really need the zero at the end, but they might want to see it). y = 10x, so 35 minutes means 350 calories, assuming it's a linear relation, which I'm pretty sure they are assuming.

OpenStudy (anonymous):

did u try it